assuming that the value of x and y are 40degree and 50degree respectively calculate for the part y

1
Expert's answer
2022-02-03T15:28:10-0500

"x+y=40\u00b0+50\u00b0=90\u00b0,"

"\frac y{x+y}=\frac {50°}{90°}=0.56."